You are driving to the grocery store at 25 m/s . You are 110 m from an intersection when the traffic light turns red. Assume that your reaction time is 0.70 s and that your car brakes with constant acceleration.

-How far are you from the intersection when you begin to apply the brakes?

-What acceleration will bring you to rest right at the intersection?

-How long does it take you to stop?

A rock is tossed straight up with a speed of 24 m/s . When it returns, it falls into a hole 15 mdeep.

-What is the rock's velocity as it hits the bottom of the hole?

-How long is the rock in the air, from the instant it is released until it hits the bottom of the hole?

You're driving down the highway late one night at17 m/s when a deer steps onto the road 44 m in front of you. Your reaction time before stepping on the brakes is 0.50 s , and the maximum deceleration of your car is 11 m/s2 .

-How much distance is between you and the deer when you come to a stop?

-What is the maximum speed you could have and still not hit the deer?

A car can go from 0 to 60 mph in 14 s .

-A second car is capable of twice the acceleration of the first car. Assuming that it could maintain the same acceleration at higher speeds, how much time will this second car take to go from 0 to 120 mph?

a11 s b3.5 s c7.0 s d14 s

In a 5.00 km race, one runner runs at a steady 12.0 km/h and another runs at 14.3 km/h .

-How long does the faster runner have to wait at the finish line to see the slower runner cross?

Chameleons catch insects with their tongues, which they can rapidly extend to great lengths. In a typical strike, the chameleon's tongue accelerates at a remarkable 240 m/s2 for 20 ms, then travels at constant speed for another 30 ms.

-During this total time of 50 ms, 1/20 of a second, how far does the tongue reach?
Answer #1

distance travelled at 25 m/s in the reaction time 0.70 sec , D1 = 25 x 0.7 = 17.5

D2 = distance travelled at constant acceleration = 110 - D1 = 110 - 17.5 = 92.5

Vo = initial speed = 25 m/s

Vf = final speed = 0 m/s

t = time taken

a = acceleration

using the equation

Vf2 = Vo2 + 2 a D2

02 = 252 + 2 a (92.5)

a = - 3.4 m/s2

Using the equation

Vf = Vo + at

0 = 25 - 3.4 t

t = 7.4 sec